OpenStudy (anonymous):

The function G(x) is described below. G(x) = 4x + 1 What is the value of G(3)? A. 16 B. 12 C. 8 D. 13

OpenStudy (jiteshmeghwal9):

since g(x)=4x+1 therefore, g(3)=4(3)+1=?

OpenStudy (anonymous):

ummm lol what?

OpenStudy (jiteshmeghwal9):

4(3)+1=?

OpenStudy (anonymous):

so g (3) is equil to 4(3) because g is 3 so 4 x 3 = what, then add one and theres your answer

OpenStudy (anonymous):

so 4 x 3 plus 1 = what?

OpenStudy (jiteshmeghwal9):

g(3) is equal to 4(3)+1, gt it ???